# Finding product sets in some classes of amenable groups

### v1

@@ 4.2. A continuous ergodic decomposition
Since $\varepsilon>0$ was arbitrary, it follows that $\lim_{N\to\infty}\frac{|(g+\Phi_{N})\cap\Phi_{N}|}{|\Phi_{N}|}=1$. Thus, $\Phi$ is a Følner sequence in $G$, and the proof is complete. ∎
6. Counterexamples on products sets
@@
To construct the counterexamples we introduce some convenient notation. We denote by $H_{3}$ the $3\times 3$ discrete Heisenberg group, that is the group of $3\times 3$ upper triangular matrices with $1$ in the diagonal and integer entries. Using the obvious Mal’cev coordinate system in this group we identify $H_{3}$ with $\mathbb{Z}^{3}$ by writing elements of $H_{3}$ as $a=(a_{1},a_{2},a_{3})$ where the group operation is given by $ab=(a_{1}+b_{1},a_{2}+b_{2},a_{3}+b_{3}+a_{1}b_{2})$. For $N\in\mathbb{N}$, we denote $[N]:=[1,N]$, where all intervals are considered in $\mathbb{Z}$, and $[N]^{\prime}:=[1,N]\cap(2\mathbb{Z}+1)$. All the counterexamples below are constructed on the group $G=H_{3}$. By Corollary 1.8, we may assume that any infinite sequence $B\subset H_{3}$ considered below is infinite in all coordinates.
@@
Then $(bc)_{1}={\rm O}(2^{Q})$ and on the other hand, $(bc)_{1}=b_{1}+c_{1}\in 2^{N}+2^{M}+[N+M]^{\prime}$, and hence we have that $(bc)_{1}={\rm O}(2^{M})$. It follows that $Q$ and $M$ are of the same order (and, in fact, we can show that they are equal). We now want to show that $t_{2}\neq 0$. By the fact that $bc\in\Phi_{Q}^{\prime}t^{-1}$, we have that
@@
Now, since $cb\in\Phi_{Q}^{\prime}t^{-1}$ and $t_{2}\neq 0$, we have that
$(bc)_{3}\gg 2^{Q}\gg 2^{M}.$
On the hand, for $M$ sufficiently large, we have that
@@ Appendix A A result of Host and Kra for amenable groups
The purpose of this appendix is to prove Lemma 3.5. The proof follows the ideas in the proof of [HK09, Proof of Proposition 6.1], adapted in our setting. We state the following classical result (see for example [Rud91, Example 11.13 (a)]), which we will need in the proof of Lemma 3.5:
@@
[HK09, Proposition 6.1 for group actions] Let $G$ be an amenable group, let $(X,\mu,T)$ be an ergodic $G$-system, $(Z,m,R)$ be its Kronecker factor and $\rho:(X,\mu,T)\to(Z,m,R)$ be a factor map. If $a\in X$ is a transitive point, then there exists a point $z\in Z$ and a Følner sequence $\Psi$ such that
@@
As in [HK09], we split the proof into two parts.
@@
where $d_{X}$ is the metric on the space $X$. Now set $z_{1}=\pi_{Z}(w_{1})$. Let $E(Z,R)$ be the Ellis semigroup of $(Z,R)$, that is, the closure of $R$ as an element of $Z^{Z}$, where this space is equipped with the pointwise convergence topology. Let $R_{0}\in\overline{(R_{h_{N}})}_{N\in\mathbb{N}}\subset E(Z,R)$. By Proposition 2.11 $R$ is a rotation, which implies that is a bijection from $Z$ to itself. In case that $(Z,m,R)$ is any distal system (and not necessarily the Kronecker factor), then we also have that $R$ is a bijection (see by [Aus88, Chapter 5]). Therefore, there exists $z_{0}\in Z$ such that $R_{0}(z_{0})=z_{1}$. Then there exists a subsequence of $(h_{N})_{N\in\mathbb{N}}$, which, by abuse of notation, we denote as $(h_{N})_{N\in\mathbb{N}}$, such that $\lim_{N\to\infty}R_{h_{N}}z_{0}=z_{1}$. Therefore, there exists a further subsequence, which once again we denote in the same way, for which it holds that
@@
$\lim_{N\to\infty}\bigg{|}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}f_{1}(T_{g}a)f_{2}(R_{g}z_{0})-\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}f_{1}(T_{g}x_{1})f_{2}(R_{g}z_{1})\bigg{|}=$

### v3

@@ 4.2. A continuous ergodic decomposition
Since $\varepsilon>0$ was arbitrary, it follows that $\lim_{N\to\infty}\frac{|(y+\Phi_{N})\cap\Phi_{N}|}{|\Phi_{N}|}=1$. Thus, $\Phi$ is a Følner sequence in $G$, and the proof of the claim is complete. ∎
Assume that $ker(s)$ is infinite, and let $(h_{n})_{n\in\mathbb{N}}$ be an enumeration of $ker(s)$ (the case when $ker(s)$ is finite is easier and can be treated with a similar argument). Let also $(E_{k})_{k\in\mathbb{N}}$ be a Følner sequence in $ker(s)$ (in case $ker(s)$ is finite, one can take $E_{k}=ker(s)$ for all $k\in\mathbb{N}$). Fix $N\in\mathbb{N}$. Then for any $n,m$ with $1\leq n,m\leq N$ and any $i,j\in\{1,\dots,\ell(N)\}$, if $g_{n}+x_{N,i}-x_{N,j}+z_{m}\in ker(s)$, then
$\lim_{k\to\infty}\frac{|(g_{n}+x_{N,i}-x_{N,j}+z_{m}+E_{k})\triangle E_{k}|}{|E_{k}|}=0.$
We can then pick $k_{N}$ large enough such that for all $n,m$ with $1\leq n,m\leq N$ and all $i,j\in\{1,\dots,\ell(N)\}$, if $g_{n}+x_{N,i}-x_{N,j}+z_{m}\in ker(s)$, then
(5.9) $\frac{|(g_{n}+x_{N,i}-x_{N,j}+z_{m}+H_{N})\triangle H_{N}|}{|H_{N}|}<\frac{1}{N},$
where we have set $H_{N}:=E_{k_{N}}$. For every $N\in\mathbb{N}$, let $\Psi_{N}=\bigsqcup_{i=1}^{\ell(N)}x_{N,i}+H_{N}$.
Claim $2$: $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$ is a Følner in $G$.
Proof of Claim $2$.
Let $y\in G$, and take $n_{y},m_{y}\in\mathbb{N}$ such that $y=g_{n_{y}}+z_{m_{y}}$. Let $\varepsilon>0$. Since $F$ is a Følner in $G$, there is $N_{0}\in\mathbb{N}$, $N_{0}>n_{y},m_{y}$, such that for every $N\geq N_{0}$,
$\frac{|(g_{n_{y}}+ker(s)+F_{N})\triangle F_{N}|}{|F_{N}|}<\frac{\varepsilon}{2}$
and $\frac{1}{N}<\frac{\varepsilon}{2}$. Let $N\geq N_{0}$. For each $i\in\{1,\dots,\ell(N)\}$, if there is $j\in\{1,\dots,\ell(N)\}$ such that $g_{n_{y}}+x_{N,i}\in x_{N,j}+ker(s)$, then this $j$ is unique by the choice of $(x_{N,i})_{i\in\{1,\dots,\ell(N)\}}$. Let
$A_{N}=\{i\in\{1,\dots,\ell(N)\}:g_{n_{y}}+x_{N,i}\in x_{N,j}+ker(s)\text{ for some }j\in\{1,\dots,\ell(N)\}\},$
and for each $i\in A_{N}$ denote the corresponding unique $j$ by $j(i)$. Let also
$B_{N}:=\{i\in\{1,\dots,\ell(N)\}:g_{n_{y}}+x_{N,i}\not\in x_{N,j}+ker(s)\text{for any }j\in\{1,\dots,\ell(N)\}\}$
and
$C_{N}:=\{j\in\{1,\dots,\ell(N)\}:g_{n_{y}}+x_{N,i}\not\in x_{N,j}+ker(s)\text{for any }i\in\{1,\dots,\ell(N)\}\},$
and observe that $\{g_{n_{y}}+x_{N,i}:i\in B_{N}\}=\{a\in G:a+ker(s)\in(g_{n_{y}}+ker(s)+F_{N})\setminus F_{N}\}$, $\{x_{N,j}:j\in C_{N}\}=\{a\in G:a+ker(s)\in F_{N}\setminus(g_{n_{y}}+ker(s)+F_{N})\}$, $|B_{N}|=|(g_{n_{y}}+ker(s)+F_{N})\setminus F_{N}|$ and $|C_{N}|=|F_{N}\setminus(g_{n_{y}}+ker(s)+F_{N})|$. We then obtain that
$(y+\Psi_{N})\triangle\Psi_{N}$ $=\Big{(}\bigsqcup_{i=1}^{\ell(N)}g_{n_{y}}+x_{N,i}+z_{m_{y}}+H_{N}\Big{)}\triangle\Big{(}\bigsqcup_{i=1}^{\ell(N)}x_{N,i}+H_{N}\Big{)}$
$\subset\Big{(}\bigcup_{i\in B_{N}}g_{n_{y}}+x_{N,i}+z_{m_{y}}+H_{N}\Big{)}\cup\Big{(}\bigcup_{i\in C_{N}}x_{N,i}+H_{N}\Big{)}\cup$
$\>\>\>\>\>\>\cup\Big{(}\bigcup_{i\in A_{N}}(g_{n_{y}}+x_{N,i}+z_{m_{y}}+H_{N})\triangle(x_{N,j(i)}+H_{N})\Big{)}.$
Therefore we have that
$|(y+\Psi_{N})\triangle\Psi_{N}|\leq|B_{N}||H_{N}|+|C_{N}||H_{N}|+\sum_{i\in A_{N}}|(g_{n_{y}}+x_{N,i}+z_{m_{y}}+H_{N})\triangle(x_{N,j(i)}+H_{N})|$
$=|(g_{n_{y}}+ker(s)+F_{N})\triangle F_{N}|||H_{N}|+\sum_{i\in A_{N}}|(g_{n_{y}}+x_{N,i}+z_{m_{y}}+H_{N})\triangle(x_{N,j(i)}+H_{N})|.$
Now for each $i\in A_{N}$, we have that $g_{n_{y}}+x_{N,i}\in x_{N,j(i)}+ker(s)$, so also $g_{n_{y}}+x_{N,i}+z_{m_{y}}\in x_{N,j(i)}+ker(s)$, and hence there is $w_{y}\in ker(s)$ such that $g_{n_{y}}+x_{N,i}+z_{m_{y}}=x_{N,j(i)}+w_{y}$. This implies $|(g_{n_{y}}+x_{N,i}+z_{m_{y}}+H_{N})\triangle(x_{N,j(i)}+H_{N})|=|x_{N,j(i)}+w_{y}+H_{N})\triangle(x_{N,j(i)}+H_{N})|=\\ |(w_{y}+H_{N})\triangle H_{N}|.$ Since $w_{y}=g_{n_{y}}+x_{N,i}-x_{N,j(i)}+z_{m_{y}}$, using (5.9) and combining with the fact that $N>n_{y},m_{y}$ we get that $\frac{|(w_{y}+H_{N})\triangle H_{N}|}{|H_{N}|}\leq\frac{1}{N}$.
Recall that $|\Psi_{N}|=\ell(N)|H_{N}|=|F_{N}||H_{N}|$ and $|A_{N}|\leq\ell(N)=|F_{N}|$, so after all we have
$\frac{|(y+\Psi_{N})\triangle\Psi_{N}|}{|\Psi_{N}|}$ $\leq\frac{|(g_{n_{y}}+ker(s)+F_{N})\triangle F_{N}||H_{N}|+\sum_{i\in A_{N}}\frac{|H_{N}|}{N}}{|F_{N}||H_{N}|}$
$\leq\frac{|(g_{n_{y}}+ker(s)+F_{N})\triangle F_{N}|}{|F_{N}|}+\frac{1}{N}<\varepsilon.$
Since $\varepsilon>0$ and $y\in G$ were arbitrary this concludes the proof of the claim. ∎
We will now prove that $\Phi,\Psi$ satisfy 1.4 for the map $s$ and therefore $G$ is square absolutely continuous. Observe that $2\Psi_{N}=\{2g:g\in\Psi_{N}\}=\{2x_{N,i}:i\in\{1,\dots,\ell(N)\}\}=\widetilde{\Phi}_{N}$ and if $u:G\to[0,1]$, then for each $N\in\mathbb{N}$ we have $\sum_{g\in\Psi_{N}}u(2g)=\sum_{g\in\bigsqcup_{i=1}^{\ell(N)}x_{N,i}+H_{N}}u(2g)=\sum_{i=1}^{\ell(N)}|H_{N}|u(2x_{N,i})=|H_{N}|\sum_{g\in\widetilde{\Phi}_{N}}u(g)$. Therefore, for each $N\in\mathbb{N}$ we have
$\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}u(2g)=\frac{1}{|\widetilde{\Phi}_{N}|}\sum_{g\in\widetilde{\Phi}_{N}}u(g)=\frac{r}{|\Phi_{N}|}\sum_{g\in\widetilde{\Phi}_{N}}u(g)\leq\frac{r}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}u(g).$
Therefore, given $\varepsilon>0$ we can take $\delta=\frac{\varepsilon}{r}$ and then for any $u:G\to[0,1]$, if
$\limsup_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}u(g)<\delta=\frac{\varepsilon}{r},$
then
$\limsup_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}u(2g)<\varepsilon.$
This concludes the proof of Theorem 1.12. ∎
6. Counterexamples on product sets
@@
To construct the counterexamples we introduce some convenient notation. We denote by $H_{3}$ the $3\times 3$ discrete Heisenberg group, that is the group of $3\times 3$ upper triangular matrices with $1$ in the diagonal and integer entries. Using the obvious Mal’cev coordinate system in this group we identify $H_{3}$ with $\mathbb{Z}^{3}$ by writing elements of $H_{3}$ as $a=(a_{1},a_{2},a_{3})$ where the group operation is given by $ab=(a_{1}+b_{1},a_{2}+b_{2},a_{3}+b_{3}+a_{1}b_{2})$. For $N\in\mathbb{N}$, we denote $[N]:=[1,N]$, where all intervals are considered in $\mathbb{Z}$, and $[N]^{\prime}:=[1,N]\cap(2\mathbb{Z}+1)$. All the counterexamples below are constructed on the group $G=H_{3}$. We may assume that any infinite sequence $B\subset H_{3}$ arising from Corollary 1.9 considered below is infinite an all coordinates.
@@
Then $2^{Q}-Q+t_{1}\leq(bc)_{1}\leq 2^{Q}+Q+t_{1}$ and on the other hand, $(bc)_{1}=b_{1}+c_{1}\in 2^{N}+2^{M}+[N+M]^{\prime}$, so $2^{N}+2^{M}-N-M\leq(bc)_{1}\leq 2^{N}+2^{M}+N+M$. Combining those with the fact that $M$ is assumed to be sufficiently large (and much larger than $N$) we obtain that $Q=M$. We now want to show that $t_{2}\neq 0$. By the fact that $bc\in\Phi_{Q}^{\prime}t^{-1}$, we have that
@@
Now, since $bc\in\Phi_{Q}^{\prime}t^{-1}$ and $t_{2}\neq 0$, we have that
$(bc)_{3}\gg 2^{Q}=2^{M}.$
On the other hand, for $M$ sufficiently large, we have that
@@ Appendix A A result of Host and Kra for amenable groups
The purpose of this appendix is to prove Lemma 3.5. The proof follows the ideas in the proof of [9, Proof of Proposition 6.1], adapted in our setting. We state the following classical result (see for example [19, Example 11.13 (a)]), which we will need in the proof of Lemma 3.5:
@@
[9, Proposition 6.1 for group actions] Let $G$ be an amenable group, let $(X,\mu,T)$ be an ergodic $G$-system, $(Z,m,R)$ be its Kronecker factor and $\rho:(X,\mu,T)\to(Z,m,R)$ be a factor map. If $a\in X$ is a transitive point, then there exists a point $z\in Z$ and a Følner sequence $\Psi$ such that
@@
As in [9], we split the proof into two parts.
@@
where $d_{X}$ is the metric on the space $X$. Now set $z_{1}=\pi_{Z}(w_{1})$. Let $E(Z,R)$ be the Ellis semigroup of $(Z,R)$, that is, the closure of $R$ as an element of $Z^{Z}$, where this space is equipped with the pointwise convergence topology. Let $R_{0}\in\overline{(R_{h_{N}})}_{N\in\mathbb{N}}\subset E(Z,R)$. By Proposition 2.12 $R$ is a rotation, which implies that is a bijection from $Z$ to itself. In case that $(Z,m,R)$ is any distal system (and not necessarily the Kronecker factor), then we also have that $R$ is a bijection (see by [3, Chapter 5]). Therefore, there exists $z_{0}\in Z$ such that $R_{0}(z_{0})=z_{1}$. Then there exists a subsequence of $(h_{N})_{N\in\mathbb{N}}$, which, by abuse of notation, we denote as $(h_{N})_{N\in\mathbb{N}}$, such that $\lim_{N\to\infty}R_{h_{N}}z_{0}=z_{1}$. Therefore, there exists a further subsequence, which once again we denote in the same way, for which it holds that
@@
$\lim_{N\to\infty}\bigg{|}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}f_{1}(T_{g}a)f_{2}(R_{g}z_{0})-\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}f_{1}(T_{g}x_{1})f_{2}(R_{g}z_{1})\bigg{|}$